OpenStudy (anonymous):

derivative of 1/(sqrt(2x^2 - 1)

OpenStudy (anonymous):

what is the derivative of \(\frac{1}{\sqrt{x}}\) ?

OpenStudy (anonymous):

x^-1

OpenStudy (anonymous):

sms can you work with me pls?

OpenStudy (anonymous):

x^-1/2

OpenStudy (anonymous):

oh yeahh

OpenStudy (anonymous):

Lets rewrite the equation in that form. easier to visualize

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[x ^{\frac{ -1 }{ 2 }}\rightarrow -\frac{ 1 }{ 2 }x ^{-\frac{ 3 }{ 2 }}\]

OpenStudy (anonymous):

same derivative rules apply. bring the exponent out front and subtract 1 from the exponent

OpenStudy (anonymous):

this problem differs in that you also need to multiply it by the dervative of what is inside the ( )

OpenStudy (anonymous):

see what you get

OpenStudy (anonymous):

helps to write \[-\frac{1}{2}x^{-\frac{3}{2}}=-\frac{1}{2\sqrt{x^3}}\]

OpenStudy (anonymous):

I would wait to rewrite it until the very end, after you have taken the derivative of the inside.

OpenStudy (anonymous):

rational exponents are helpful when finding derivatives, but you need to figure out what you really get at the end. now you can use the chain rule. since you see the derivative of \[\frac{1}{\sqrt{x}}\] is \[-\frac{1}{2\sqrt{x^2}}\] you know that the derivative of \[\frac{1}{\sqrt{f(x)}}\] is \[-\frac{f'(x)}{2\sqrt{f^3(x)}}\]

OpenStudy (anonymous):

How are you doing Jay

OpenStudy (anonymous):

the derivative of the inside is 4x^3 right?

OpenStudy (anonymous):

or you could do whats @smstevns said if it makes it easier for you

OpenStudy (anonymous):

remember it is minus 1

OpenStudy (anonymous):

no it is \(4x\)

OpenStudy (anonymous):

satellite proved the theorem for you to remember in the future

OpenStudy (anonymous):

this is what you need to write \[-\frac{f'(x)}{2\sqrt{f^3(x)}}\] replace \(f(x)\) by \(2x^2-1\) and replace \(f'(x)\) by \(4x\)

OpenStudy (anonymous):

It can be simplified from there depending on what your teacher is looking for. -4x/2 = -2x

OpenStudy (anonymous):

simplifying the numerator

OpenStudy (anonymous):

i see sorry

OpenStudy (anonymous):

Hope we helped. peace